# GraphTranslator: Aligning Graph Model to Large Language Model for Open-ended Tasks

### 11 Feb 2024 v1

@@ Abstract.
Large language models (LLMs) like ChatGPT, exhibit powerful zero-shot and instruction-following capabilities, have catalyzed a revolutionary transformation across diverse research fields of artificial intelligence, especially for open-ended tasks. While the idea is less explored in the graph domain, despite the availability of numerous powerful graph models (GMs), they are restricted to tasks in a pre-defined form. Although several methods applying LLMs to graphs have been proposed, they fail to simultaneously handle the pre-defined and open-ended tasks, with LLM as a node feature enhancer or as a standalone predictor. To break this dilemma, we propose to bridge the pretrained GM and LLM by a Translator, named GraphTranslator, aiming to leverage GM to handle the pre-defined tasks effectively and utilize the extended interface of LLMs to offer various open-ended tasks for GM. To train such Translator, we propose a Producer capable of constructing the graph-text alignment data along node information, neighbor information and model information. By treating the node representation as a type of language, the proposed GraphTranslator empowers an LLM to make predictions based on node representation and language instructions, providing a unified perspective for both pre-defined and open-ended tasks. Extensive results show that the proposed GraphTranslator effectively improves the results of zero-shot node classification. The graph question answering experiments reveal our GraphTranslator potential across a broad spectrum of open-ended applications through language instructions.
@@
copyright: acmcopyrightjournalyear: 2024conference: Proceedings of the ACM Web Conference 2024; May 13–17, 2024; Singaporeccs: Computer systems organization Embedded systemsccs: Networks Network reliability
@@ 1. Introduction
Recently, the emergence of large language models (LLMs) like ChatGPT  [^1] https://openai.com/blog/chatgpt has brought about a paradigm shift in natural language processing (NLP) research, showcasing their impressive emergent abilities (Wei et al., 2022b) for flexible open-ended tasks based on natural language instructions. Such development of LLMs also has revolutionized research of diverse modality (Li et al., 2023; Wu et al., 2023; Liu et al., 2023a). Taking image as an example, LLMs largely facilitates the visual-centric open-ended tasks, such as instructed image-to-text generation and visual question answering tasks, transforming how we interpret and interact with visual information.
@@
In this paper, to address above challenges, we propose a novel framework called GraphTranslator to align the pre-trained GM to LLM, solving both pre-defined and open-ended tasks. Specifically, for the challenge of modality gap, GraphTranslator introduces a Translator module which to converts node embedding into token embedding space, which can be interpreted by LLM. To achieve this, the Translator module learns a set of graph queries to extract the language information of node embeddings, then performs generative learning for adapting to LLM. For the second challenge of lacking alignment data, we introduce a Producer that capable of constructing (node embedding, textual description) pairs through the powerful generation ability of LLMs. To seamlessly textualize the information encoded in node embeddings, we generate the description step by step, including briefing node’s attribute, summarizing neighbor attribute, then reasoning their commonality. After training on the alignment data, as presented in Figure 1 (b), a frozen LLM, equipped with Translator, can handle various open-ended tasks based on the node embedding and instructions. In conclusion, in our GraphTranslator , pre-defined tasks can be tackled efficiently by the customized GM, while LLMs further provide GM with interactivity and interpretability.
@@
  • We evaluate our method on real-world datasets for open-ended tasks. The experimental results demonstrate the effectiveness of GraphTranslator on zero-shot node classification. The graph question answering experiments reveal the noteworthy potential of GraphTranslator when applied to tasks predicated upon language instructions.
@@ 2.1. Notations and Problem Settings
Open-ended Tasks  On the contrary, open-ended tasks offer greater flexibility and freedom, characterized by the absence of explicit task specifications or evaluation criteria. Models designed for open-ended tasks often depend on autonomous learning and creative problem-solving approaches. In real-world scenarios, new tasks often emerge with evolving business requirements, such as classifying new labels or tasks driven entirely by human instructions. The computer vision community has also adopted the language instruction paradigm for tasks like image-to-text generation (Liu et al., 2023a; Li et al., 2023). However, current graph models are constrained by predefined tasks and lack the flexibility to accommodate open-ended task customization guided by language instructions like LLMs.
@@ 2.2. Overall Architecture
The primary goal of our proposed GraphTranslator is to align graph models to LLMs, in order to harness the emergent capabilities of LLMs for open-ended tasks. Specifically, GraphTranslator consists of four components: (1) Frozen graph model (GM), is pre-trained on a large-scale graph, such as e-commerce graphs with billions of nodes, yielding embeddings for all nodes that encoding the graph’s information for downstream tasks. We use the pre-trained GraphSAGE model as an example in this work. (2) Frozen LLM, is trained on broad text corpus, showcasing emergent abilities when the number of parameters reach a certain scale. We employ the pre-trained ChatGLM2-6B for demonstration. (3) Producer, is designed to construct the alignment data for training the Translator module, i.e., (node embedding, textual description) pairs. (4) Translator module, is designed to project the GM learned node embeddings into LLM space, eliminating the modality gap between GM and LLM.
@@ 2.3. Frozen Graph Model
The representation learning of TAGs has been extensively studied. Given a TAG $\mathcal{G}=\left(\mathcal{V},\boldsymbol{A},\left\{s_{v}\right\}_{v\in\mathcal{V}}\right)$, the typical graph neural networks (GNNs) (Kipf and Welling, 2016; Velickovic et al., 2017; Hamilton et al., 2017) is denoted as $g_{\boldsymbol{\theta}}(\boldsymbol{A},\boldsymbol{X})$, where $\boldsymbol{\theta}$ is set of learnable parameters, $\boldsymbol{X}$ is the node features processed by shallow methods such as bag-of-words (BoW) (Harris, 1954) or skip-gram (Mikolov et al., 2013). Taking the GraphSAGE (Hamilton et al., 2017) as an example, typically, GraphSAGE samples a fixed-size neighbors $\mathcal{N}(v)$ around target node $v$, then concatenate the node’s previous layer embedding $\boldsymbol{h}_{v}^{k-1}$ with the aggregated neighborhood vectors $\{\boldsymbol{h}^{k-1}_{u},\forall u\in\mathcal{N}(v)\}$ by:
@@ 2.4. Frozen Large Language Model
LLMs are trained on extensive text corpora, acquiring a substantial amount of knowledge. It’s worth noting that LLMs reveal their capabilities only when they reach a certain parameter scale (Wei et al., 2022a). To prevent the potential issues of catastrophic forgetting and the unbearable training costs associated with handling a large number of parameters, here we keep the LLM parameters fixed. Specifically, we employ ChatGLM2-6B, which is open-source bilingual (Chinese-English) language model. ChatGLM2-6B employs a specific type of autoregressive blank infilling task, which aligns with the typical design philosophy of most pretraining tasks. This approach involves a ”disrupt and reconstruct” strategy, wherein portions of the original text are masked (disrupted) and subsequently predicted (reconstructed). After extensive training on large-scale corpora, these LLMs acquires the ability to retain a considerable amount of knowledge and provide reasonable answers to human queries.
@@ 2.5. Producer Module
To align GM and LLM, we construct the alignment data $P=\{\boldsymbol{z}_{v},t_{v}\}^{N_{P}}_{i=1}$, where $t_{v}$ outline the information encoded within node embedding $\boldsymbol{z}_{v}$ for each node $v\in\mathcal{V}_{P}$ This process is not only related to graph data but also intertwined with the design of the GMs, so we employ LLM to construct high-quality description text with Chain-of-Thought (CoT). Taking GraphSAGE as an example, we have devised a pipeline that guides LLM in constructing descriptive information along three key dimensions:
$\bullet$ Node Information: Node information of node is highly significant and is preserved in node embedding. Generally, node attributes, such as text or numerical data, are considered as one of the features for each node. These attributes are transformed into node features, which is achieved by Bag-of-Word or word embeddings models. Therefore, the Producer uses LLM to summarize and analyze the attributes of each node $v$ in the training set, yielding the self information description, denoted as $t^{s}_{v}$.
$\bullet$ Neighbor Information: GraphSAGE also consider neighboring information. GraphSAGE randomly samples a subset of neighboring nodes $\mathcal{N}(v)$ and aggregate their representations, yielding the neighbor embedding. Node self and neighbor information are further fused through weighted summation or concatenation. The Producer proceeds to employ LLM to summarize the attributes of the sampled neighbors $\mathcal{N}(v)$, resulting in the neighbor information description, denoted as $t_{v}^{\mathcal{N}(v)}$.
$\bullet$ Model Bias: Given that most GNNs tend to uncover similarities between nodes and their neighbors for smoothing purposes, we instruct LLM to consolidate shared information. Therefore, the Producer further utilizes LLM to infer the commonalities between node $v$ and its neighbors $\mathcal{N}(v)$ based on $t_{v}^{s}$ and $t_{v}^{\mathcal{N}(v)}$, resulting in commonality information denoted as $t_{v}^{c}$.
@@ 2.6. Translator Module
There exists modality gap between the trained GM and LLM, LLMs fail to interpret node representations. Namely, the sizes of the node embedding and the input token of LLM are different, and they have different feature space. To resolve this discrepancy, inspired by (Li et al., 2023), we introduce the Translator module, which aims to align GM and LLM by converting the learned node embedding into token representations. A naive solution is applying a simple trainable projection matrix can convert $\boldsymbol{z}_{\boldsymbol{v}}$ into language embedding tokens, aligning their dimensionality with that of the word embedding space within the language model. While the simple transformation is hard to extract and translate the complex information contained within node representations to natural language, and may struggle to generalize to unseen nodes.
In our Translator module, for a pair $(\boldsymbol{z}_{v},t_{v})$ in alignment data, we utilize two encoders, denoted $f_{z}(\cdot)$ and $f_{t}(\cdot)$, to extract their language features for alignment. For textual description $t_{v}$, we leverage the text encoder $f_{t}(t_{v})$ (e.g., BERT (Devlin et al., 2018)) to extract the language features $\boldsymbol{T}_{v}=f_{t}(t_{v})$, where $f_{t}(\cdot)$ contains 12 layers of Transformer blocks. For node embedding $\boldsymbol{z}_{v}$, we also adopt a transformer-based network $f_{z}(\cdot)$ with $M$ learnable token embeddings as input, termed query tokens $\boldsymbol{Q}=\left\{\boldsymbol{q}_{i}\right\}_{i=1}^{M}$, and output $M$ features $\boldsymbol{H}_{v}=\{\boldsymbol{h}_{v,i}\}^{M}_{i=1}$ and $\boldsymbol{H}_{v}=f_{z}(\boldsymbol{Q},\boldsymbol{z}_{v})$, extracting the information of $\boldsymbol{z}_{v}$ that is most related to $t_{v}$. To achieve this, as inspired by (Li et al., 2023), the query tokens $\boldsymbol{Q}$ are designed to interact with each another using self-attention layers, interface with node embedding $\boldsymbol{z}_{v}$ through cross-attention layers, and communicate with description $t_{v}$ by harmonizing the self-attention layers between $f_{t}$ and $f_{v}$.
@@ 2.7. Model Training
Inspired by (Li et al., 2023), we train the lightweight Translator module following a two-stage joint training paradigm, bridging the gap between graph and LLMs step by step. In the first stage, we train the Translator module for extracting $\boldsymbol{H}_{v}$ from the node embedding $\boldsymbol{z}_{v}$ most relevant to $t_{v}$. In the second stage, we perform the generative learning by connecting the output of the Translator to the frozen LLM. We continue to training the Translator such that its output can be understood by the LLM.
@@ 3.1. Experimental Setting
$\bullet$ ArXiv dataset is a graph constructed by a collection of ArXiv research papers with 169,343 nodes and 1,166,243 edges. Each node represents an individual research paper with textual attributes including the paper’s title and abstract, and the edges reflect the paper citation relationship. For training the graph model, we divide nodes into 90,941 training nodes and 29,799 validation nodes following (Hu et al., 2020). In consideration of ChatGLM2-6B’s inference speed, our test set contains 4,000 data points for 40 computer science categories, chosen proportionally based on the labels in the public split.
@@
We compare our model with several pre-trained transformer-based language models. In the zero-shot scenario, we calculate the similarity between the condensed node description and the labeled text with a prompt, and predict the most similar class.
@@
$\bullet$ RoBERTa (Liu et al., 2019a) and $*{}^{*}$: RoBERTa is a variant of the BERT model which incorporates additional training techniques and data augmentation strategies to improve the model’s performance further. Similar to $*{}^{*}$, we finetune RoBERTa to obtain $*{}^{*}$. To further examine the effectiveness of our GraphTranslator  aligning GM to LLM, we compare with the methods only using text to query LLM:
@@
Graph model. In the Taobao dataset, we utilize GraphSAGE (Hamilton et al., 2017) to aggregate neighbors’ information and generate node embeddings. The GraphSAGE model employs a 2-layer aggregation and samples 10 neighbors for each layer. The dimensions of the input, intermediate, and output layers are set to 768. For the ArXiv dataset, we configure the intermediate dimension as 1024, while maintaining the remaining settings identical to the Taobao dataset.

### 28 Feb 2024 v4

@@ Abstract.
Large language models (LLMs) like ChatGPT, exhibit powerful zero-shot and instruction-following capabilities, have catalyzed a revolutionary transformation across diverse fields, especially for open-ended tasks. While the idea is less explored in the graph domain, despite the availability of numerous powerful graph models (GMs), they are restricted to tasks in a pre-defined form. Although several methods applying LLMs to graphs have been proposed, they fail to simultaneously handle the pre-defined and open-ended tasks, with LLM as a node feature enhancer or as a standalone predictor. To break this dilemma, we propose to bridge the pretrained GM and LLM by a Translator, named GraphTranslator, aiming to leverage GM to handle the pre-defined tasks effectively and utilize the extended interface of LLMs to offer various open-ended tasks for GM. To train such Translator, we propose a Producer capable of constructing the graph-text alignment data along node information, neighbor information and model information. By translating node representation into tokens, GraphTranslator empowers an LLM to make predictions based on language instructions, providing a unified perspective for both pre-defined and open-ended tasks. Extensive results demonstrate the effectiveness of our proposed GraphTranslator on zero-shot node classification. The graph question answering experiments reveal our GraphTranslator potential across a broad spectrum of open-ended tasks through language instructions. Our code is available at: https://github.com/alibaba/GraphTranslator.
@@
journalyear: 2024copyright: acmlicensedconference: Proceedings of the ACM Web Conference 2024; May 13–17, 2024; Singapore, Singaporebooktitle: Proceedings of the ACM Web Conference 2024 (WWW ’24), May 13–17, 2024, Singapore, Singaporedoi: 10.1145/3589334.3645682isbn: 979-8-4007-0171-9/24/05ccs: Computing methodologies Natural language generationccs: Information systems Wrappers (data mining)
@@ 1. Introduction
Recently, the emergence of large language models (LLMs) like ChatGPT  [^1] https://openai.com/blog/chatgpt has brought about a paradigm shift in natural language processing (NLP) research, showcasing their impressive emergent abilities (Wei et al., 2022a) for flexible open-ended tasks based on natural language instructions. Such development of LLMs also has revolutionized research of diverse modality (Li et al., 2023; Wu et al., 2023; Liu et al., 2023a). Taking image as an example, LLMs largely facilitates the visual-centric open-ended tasks, such as instructed image-to-text generation and visual question answering tasks, transforming how we interpret and interact with visual information.
@@
In this paper, to address above challenges, we propose a novel framework called GraphTranslator to align the pre-trained GM to LLM, solving both pre-defined and open-ended tasks. Specifically, for the challenge of modality gap, GraphTranslator introduces a Translator module which converts node embedding into token embedding space. To achieve this, the Translator module learns a set of graph queries to extract the language information of node embeddings, then performs generative learning for adapting to LLM. For the second challenge of lacking alignment data, we introduce a Producer that capable of constructing (node embedding, textual description) pairs through the powerful generation ability of LLMs. To seamlessly textualize the information encoded in node embeddings, we generate the description step by step, including briefing node’s attribute, summarizing neighbor attribute, then reasoning their commonality. After training on the alignment data, as presented in Figure 1 (b), a frozen LLM, equipped with Translator, can handle various open-ended tasks based on the node embedding and instructions. In conclusion, in our GraphTranslator , pre-defined tasks can be tackled efficiently by the customized GM, while LLMs further provide GM with interactivity and interpretability.
@@
  • The experimental results on real-world datasets demonstrate the effectiveness of GraphTranslator on zero-shot node classification. The graph question answering experiments reveal the noteworthy potential of GraphTranslator when applied to tasks predicated upon language instructions.
@@ 2.1. Notations and Problem Settings
Open-ended Tasks  On the contrary, open-ended tasks offer greater flexibility, characterized by the absence of explicit task specifications or evaluation criteria. Models designed for open-ended tasks often depend on autonomous learning and creative problem-solving approaches. In real-world scenarios, new tasks often emerge with evolving business requirements, such as classifying new labels or tasks driven entirely by human instructions. The computer vision community has also adopted the language instruction paradigm for tasks like image-to-text generation (Liu et al., 2023a; Li et al., 2023). However, current graph models are constrained by predefined tasks and fail to adopt to open-ended task customization guided by language instructions like LLMs.
@@ 2.2. Overall Architecture
The primary goal of our GraphTranslator is to align graph models to LLMs, to leverage the emergent capabilities of LLMs for open-ended tasks. Specifically, GraphTranslator consists of four components: (1) Frozen graph model (GM), is pre-trained on a large-scale graph, such as e-commerce graphs with billions of nodes, yielding embeddings for all nodes that encoding the graph’s information for downstream tasks. We use the pre-trained GraphSAGE model as an example in this work. (2) Frozen LLM, is trained on broad text corpus, showcasing emergent abilities when the number of parameters reach a certain scale. We employ the pre-trained ChatGLM2-6B for demonstration. (3) Producer, is designed to construct the alignment data for training the Translator module, i.e., (node embedding, textual description) pairs. (4) Translator module, is designed to project the GM learned node embeddings into LLM space, eliminating the modality gap between GM and LLM.
@@ 2.3. Frozen Graph Model
The representation learning of TAGs has been extensively studied. Given a TAG $\mathcal{G}=\left(\mathcal{V},\boldsymbol{A},\left\{s_{v}\right\}_{v\in\mathcal{V}}\right)$, the typical graph neural networks (GNNs) (Kipf and Welling, 2016; Velickovic et al., 2017; Hamilton et al., 2017) is denoted as $g_{\boldsymbol{\theta}}(\boldsymbol{A},\boldsymbol{X})$, where $\boldsymbol{\theta}$ is set of learnable parameters, $\boldsymbol{X}$ is the node features processed by shallow methods such as bag-of-words (BoW) (Harris, 1954) or skip-gram (Mikolov et al., 2013). Taking the GraphSAGE (Hamilton et al., 2017) as an example, typically, GraphSAGE samples a fixed-size neighbors $\mathcal{N}(v)$ around target node $v$, then concatenates the node’s previous layer embedding $\boldsymbol{h}_{v}^{k-1}$ with the aggregated neighborhood vectors $\{\boldsymbol{h}^{k-1}_{u},\forall u\in\mathcal{N}(v)\}$ by:
@@ 2.4. Frozen Large Language Model
LLMs are trained on extensive text corpora, acquiring a substantial amount of knowledge. It’s worth noting that LLMs reveal their capabilities only when they reach a certain parameter scale (Wei et al., 2022a). To prevent the potential issues of catastrophic forgetting and the unbearable training costs associated with handling a large number of parameters, here we keep the LLM parameters fixed. Specifically, we employ ChatGLM2-6B, which is open-source bilingual (Chinese-English) language model. ChatGLM2-6B employs a specific type of autoregressive blank infilling task, which aligns with the typical design philosophy of most pretraining tasks. This approach involves a ”disrupt and reconstruct” strategy, wherein portions of the original text are masked (disrupted) and subsequently predicted (reconstructed). After extensive training on large-scale corpora, these LLMs acquire the ability to retain a considerable amount of knowledge and provide reasonable answers to human queries.
@@ 2.5. Producer Module
To align GM and LLM, we construct the alignment data $P=\{(\boldsymbol{z}_{v},t_{v})\}^{N_{P}}_{i=1}$, where $t_{v}$ outline the information encoded within node embedding $\boldsymbol{z}_{v}$ for each node $v\in\mathcal{V}_{P}$ This process is not only related to graph data but also intertwined with the design of the GMs, so we employ LLM to construct high-quality description text with Chain-of-Thought (CoT). Taking GraphSAGE as an example, we have devised a pipeline that guides LLM in constructing descriptive information along three key dimensions:
$\bullet$ Node Information: Generally, node attributes, such as text or numerical data, are considered as the features for each node. In GMs, these attributes are often transformed into node features, which is achieved by Bag-of-Word or word embeddings models. Therefore, the Producer uses LLM to summarize and analyze the attributes of each node $v$ in the training set, yielding node description, denoted as $t^{s}_{v}$.
$\bullet$ Neighbor Information: GraphSAGE also consider neighboring information. GraphSAGE randomly samples a subset of neighboring nodes $\mathcal{N}(v)$ and aggregate their representations, yielding the neighbor embedding. Node and neighbor information are further fused through weighted summation or concatenation. So our Producer employs LLM to summarize the attributes of the sampled neighbors $\mathcal{N}(v)$, resulting in the neighbor information description, denoted as $t_{v}^{\mathcal{N}(v)}$.
$\bullet$ Model Information: Given that most GNNs tend to uncover similarities between nodes and their neighbors for smoothing purposes, we instruct LLM to summarize the shared information. Therefore, the Producer further utilizes LLM to infer the commonalities between node $v$ and its neighbors $\mathcal{N}(v)$ based on $t_{v}^{s}$ and $t_{v}^{\mathcal{N}(v)}$, resulting in commonality information denoted as $t_{v}^{c}$.
@@ 2.6. Translator Module
There exists modality gap between the trained GMs and LLMs, so LLMs fail to interpret node representations. Namely, the sizes of the node embedding and the input token of LLM are different, and they have different feature space. To resolve this discrepancy, inspired by (Li et al., 2023), we introduce the Translator module, which aims to align GM and LLM by converting the learned node embedding into token representations. A naive solution is applying a simple trainable projection matrix can convert $\boldsymbol{z}_{\boldsymbol{v}}$ into language embedding tokens, aligning their dimensionality with that of the word embedding space within the language model. While the simple transformation is hard to extract and translate the complex information contained within node representations to natural language, and may struggle to generalize to unseen nodes.
In our Translator module, for a pair $(\boldsymbol{z}_{v},t_{v})$ in alignment data, we utilize two encoders, denoted as $f_{z}(\cdot)$ and $f_{t}(\cdot)$, to extract their language features for alignment. For textual description $t_{v}$, we leverage the text encoder $f_{t}(t_{v})$ (e.g., BERT (Devlin et al., 2018)) to extract the language features $\boldsymbol{T}_{v}=f_{t}(t_{v})$, where $f_{t}(\cdot)$ contains 12 layers of Transformer blocks. For node embedding $\boldsymbol{z}_{v}$, we also adopt a transformer-based network $f_{z}(\cdot)$ with $M$ learnable token embeddings as input, termed query tokens $\boldsymbol{Q}=\left\{\boldsymbol{q}_{i}\right\}_{i=1}^{M}$, and output $M$ features $\boldsymbol{H}_{v}=\{\boldsymbol{h}_{v,i}\}^{M}_{i=1}$ and $\boldsymbol{H}_{v}=f_{z}(\boldsymbol{Q},\boldsymbol{z}_{v})$, extracting the information of $\boldsymbol{z}_{v}$ that is most related to $t_{v}$. To achieve this, as inspired by (Li et al., 2023), the query tokens $\boldsymbol{Q}$ are designed to interact with each another using self-attention layers, interface with node embedding $\boldsymbol{z}_{v}$ through cross-attention layers, and communicate with description $t_{v}$ by sharing the self-attention layers between $f_{t}$ and $f_{v}$.
@@ 2.7. Model Training
Inspired by (Li et al., 2023), we train the lightweight Translator module following a two-stage training paradigm, bridging the gap between graph and LLMs step by step. In the first stage, we train the Translator module for extracting $\boldsymbol{H}_{v}$ from the node embedding $\boldsymbol{z}_{v}$ most relevant to $t_{v}$. In the second stage, we perform the generative learning by connecting the output of the Translator to the frozen LLM. We continue to training the Translator such that its output can be understood by LLM.
@@ 3.1. Experimental Setting
$\bullet$ ArXiv dataset is a graph constructed by a collection of ArXiv research papers with 169,343 nodes and 1,166,243 edges. Each node represents an individual research paper with textual attributes including the paper’s title and abstract, and the edges reflect the paper citation relationship. For training the graph model, we divide nodes into 90,941 training nodes and 29,799 validation nodes following (Hu et al., 2020). In consideration of LLM’s inference speed, our test set contains 4,000 data points for 40 computer science categories, chosen proportionally based on the labels in the public split.
@@
We compare our model with several pre-trained transformer-based language models. In the zero-shot scenario, we calculate the similarity between node description and the labeled text with a prompt, and predict the most similar class.
@@
$\bullet$ RoBERTa (Liu et al., 2019a) and $*{}^{*}$: RoBERTa is a variant of the BERT model which incorporates additional training techniques and data augmentation strategies to improve the model’s performance further. Similar to $*{}^{*}$, we finetune RoBERTa to obtain $*{}^{*}$.
To further examine the effectiveness of our GraphTranslator , we compare with the methods only using text to query LLM:
@@
Graph Model. In Taobao dataset, we utilize GraphSAGE (Hamilton et al., 2017) to generate node embeddings. The GraphSAGE model employs a 2-layer aggregation and samples 10 neighbors for each layer. The dimensions of the intermediate layer isset to 768. For the ArXiv dataset, we configure the intermediate dimension as 1024, while maintaining the remaining settings identical to the Taobao dataset.